Accounting Cycle Review E 0-4 A tabular analysis of the transactions made during August 2017 by Cullumber Company during its first month of operations is shown below. Each increase and decrease in stockholders’ equity is explained. Assets = Liabilities + Stockholders’ Equity Cash + A/R + Supp. + Equip. = Accounts Payable + Common Stock + Retained Earnings Revenues – Expenses – Dividends (1) $21,400 $21,400 Com. Stock (2) –1,500 $5,100 $3,600 (3) –800 $800 (4) 4,100 $6,400 $10,500 Serv. Rev. (5) –2,000 –2,000 (6) –1,600 –$1,600 Div. (7) –850 –$850 Rent Exp. (8) 530 –530 (9) –2,500 –2,500 Salar. Exp. (10) 320 –320 Util. Exp. (b) Determine how much stockholders’ equity increased for the month. Increase in stockholders’ equity $ (c) Compute the net income for the month.
Answer #1
S.No. Assets Liabilities Stockholder's Equity
Cash+ A/R+ Supplies+ Equipment= Accounts Payable+ Common Stock+ Retained Earnings
Revenues- Expenses- Dividend
1 $                                                                    21,400.00 $          21,400.00 Common Stock
2 $                                                                     -1,500.00 $     5,100.00 $                3,600.00
3 $                                                                         -800.00 $       800.00
4 $                                                                       4,100.00 $   6,400.00 $ 10,500.00 Service Revenue
5 $                                                                     -2,000.00 $               -2,000.00
6 $                                                                     -1,600.00 -1600 Dividend
7 $                                                                         -850.00 $     -850.00 Rent Expenses
8 $                                                                          530.00 $     -530.00
9 $                                                                     -2,500.00 $ -2,500.00 Salary Expenses
10 $                                                                         -320.00 $     -320.00 Utility Expenses
Cullumber Company
Stockholder's Equity
Increase in Common Stock=(A) $ 21,400.00
Add: Service Revenue=(B) $ 10,500.00
Less:
Expenses
Dividend $   1,600.00
Rent Expenses $       850.00
Salary Expenses $   2,500.00
Utility Expenses $       320.00
Total Expenses=(C ) $   5,270.00
Increase in stockholdrer's Equity=(A)+(B)-(C ) $ 26,630.00
Net Income
Revenue
Service Revenue=(A) $ 10,500.00
Expenses
Rent Expenses $       850.00
Salary Expenses $   2,500.00
Utility Expenses $       320.00
Total Expenses=(B) $   3,670.00
Net Income=(A)-(B) $   6,830.00